OpenStudy (anonymous):

solve: -1/2|x+7|=6 i got (-19,5) is this correct if not can you show me step by step

OpenStudy (anonymous):

is it a fraction at the left?

OpenStudy (anonymous):

It's right :)

OpenStudy (anonymous):

i don't think there should even be answers for this...

OpenStudy (anonymous):

if you multiply both sides by -2 you get an absolute value equal to a negative.

OpenStudy (phi):

if that is =, then you will never get it to work

OpenStudy (anonymous):

multiply both sides by -2 you get |x+7|=-12

OpenStudy (anonymous):

yeah a fraction at the left

OpenStudy (anonymous):

it's gonna be like x+7=-12 or x+7=12 then solve for x

OpenStudy (anonymous):

yeah i did that

OpenStudy (anonymous):

Yeah, but try it now. try substituting -19 or 5 at the original equation, you will see it won't work out.

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

because absolute value is never negative, and your equation simplifies to |x+7| = -12

OpenStudy (anonymous):

thats what i was forgetting then was the absolute sign. for some reason i toook it out idk why lol

OpenStudy (phi):

with these absolute value problems, you always have to check for "spurious answers"

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks all